A sailboat under auxiliary power is proceeding on a bearing 25° north of west at 6,25km/h in still water.Then a tail wind blowing 15km/h in the direction 35° south of west, alters the course of the sailboat.What is the resultant speed and direction of the saiboat (correct to 2 decimal places)?
1
Expert's answer
2020-10-17T16:33:39-0400

"\\textsf{Calculuating\u00ad the components of vector}\\\\ A_1 A_2, A_3 A_2 \\,\\textsf{and}\\, A_3 A_1\\,\\textsf{in the system of rectangular axes}\\\\\\textsf{and}\\,A_2\\, \\textsf{is a point}\\, 6.25\\,\\textsf{units away from}\\\\ A \\,\\textsf{at a bearing of}\\, 25\\degree, A_3 \\,\\textsf{is a point}\\\\ 15 \\,\\textsf{units away from}\\, A \\\\\\textsf{at a bearing of}\\, 215\\degree \\\\\n\n\n\\textsf{We first construct a diagram including}\\\\\\textsf{\u00adpoint}\\, A_1, A_2, A_3 \\,\\textsf{at the given bearing}\\\\\\textsf{an\u00add determine the magnitude of}\\, A_3 A_2. \\\\\\textsf{Since the magnitude of velocity}\\\\\\textsf{i\u00ads speed (see the figure).}\\\\\n\n\n\\textsf{Hence, the}\\, x \\, \\textsf{and}\\, y \\,\\textsf{components\u00ad of vector}\\\\ \\overrightarrow{A_1 A_2},\u00ad \\,\\textsf{and}\\,\\overrightarrow{A_3 A_1}\\, \\textsf{are respectively given by} \\\\\n\n{A_1 A_2}_x = -6.25\\cos(25\\degree)\u00ad = -5.66\\\\\n\n{A_1 A_2}_y = 6 \\sin(25\\degree) = 2.64\\\\\n\n\\overrightarrow{A_1 A_2}\\,\\textsf{can be written as}\\\\\n\n\\overrightarrow{A_1 A_2\u00ad} = {A_1 A_2}_x \\textbf{i} + {A_1 A_2}_y\\textbf{j}, \\\\\\textsf{where}\\, \\textbf{i} \\,\\textsf{and}\\, \\textbf{j} \\,\\textsf{are unit vectors.}\\\\\n\n\\therefore \\overrightarrow{A_1 A_2} = -5.66\\textbf{i} + 2.64\\textbf{j}\\\\\n\n\n\\textsf{Similarly,} \\\\\n\n{A_3 A_1}_x = 15\\cos(215\\degree) = -12.29\\\\\n\n{A_3 A_1}_y = 15 \\sin(215\\degree) = -8.60\\\\\n\n\\overrightarrow{A_3 A_1} = -12.29\\textbf{i} - 8.60\\textbf{j}\\\\\n\n\n\\textsf{By the vector law of addition,}\\\\\n\n\\overrightarrow{A_3 A_1} + \\overrightarrow{A_1 A_2} = \\overrightarrow{A_3 A_2} \\\\\n\n\\therefore \\overrightarrow{A_3 A_2} = -17.95\\textbf{i} - 5.96\\textbf{j}\\\\\n\n|A_3 A_2| \\, \\textsf{is the resultant speed and}\\\\\\textsf{direct\u00adion of the sail boat.}\\\\\n\n\n\\therefore |A_3 A_2| = \\sqrt{17.95^2 + 5.96^2} = 18.91\\\\\n\n\n\\textsf{Therefore, the resultant speed is}\\, 18.91 \\,\\textit{km\/\u00adh,}\\,\\textsf{to}\\,2\\textsf{d.p.}\\\\\\textsf{Also, the direction of the sail boat is given by}\\\\\n\n\\theta = \\arctan\\left(\\frac{5.96}{17.95}\\right) = 18.37\\degree\\\\\n\n\n\n\\therefore \\textsf{\nThe direction of the sail boat is}\\, W\\,18.37\\degree S"